# Fast wavelet basis search for generic gravitational wave bursts in Pulsar Timing Array data

### v1

@@ II.1 Models
where $\bm{H}_{+}$ and $\bm{H}_{\times}$ are the plus and cross GW polarizations respectively, and have independent amplitudes $A_{+}$ and $A_{\times}$, with independent phases $\phi_{0,+}$ and $\phi_{0,\times}$. This leaves $t_{0},f_{0}$, and $\tau$ as common parameters between both polarizations.
@@
where $F_{+}(\Omega_{k},\Omega_{\rm gw})$ and $F_{\times}(\Omega_{k},\Omega_{\rm gw})$ are the plus and cross polarization mode antenna responses, respectively. For further details on the geometric response, see [25]. Eq. (10) assumes that the only contribution to the signal is from the Earth term (ET). This is a result of our current observing baseline ($\sim 20$ years) being significantly shorter than the light travel time between the Earth and any pulsar in the PTA ($\sim 10^{3}yrs$). Thus, any GW bursts which appear in only one pulsar will not affect the residuals in any other pulsar, nor will it reach the Earth within the time frame of the experiment. As such, this burst will appear as a single noise transient in that particular pulsar.
Moreover, it is expected that the number of events that we can possibly observe in the ET only search scales with the number of pulsars $\sim N_{p}^{3/2}$. However, the pulsar term (PT) only search scales as $\sim N_{p}$ due to the extended timing baseline when combining individual pulsar data. As a result, the number of possible observed events in the ET only search exceeds the number of possible observed events in the PT only search by $\sim N_{p}^{1/2}$. Considering current PTAs contain of order 100 pulsars, there’s potential to observe 10$\sim$ 10$\times$ 10 × more events in the ET only search. For additional details on these arguments, see [7, 26].
@@ II.2 Likelihood
$\Sigma=B^{-1}+T^{T}N^{-1}T.$ (14)
B is the prior matrix for the hyper-parameters, and T is the design matrix for the timing model, red noise (RN) and jitter (see [28]).
@@
$\log L=\sum_{i=1}^{N_{\mathrm{psr}}}\left[-\frac{1}{2}(\delta\bm{t}_{i}|\delta\bm{t}_{i})-\frac{1}{2}\log{\det(2\pi C_{i})}+\sum_{k=1,2}\sigma_{ki}N_{i}^{k}-\frac{1}{2}\sum_{k=1,2}\sum_{l=1,2}\sigma_{ki}\sigma_{li}M_{i}^{kl}\right],$ (17)
where the the last two terms are inner products between the $i^{th}$ pulsar residuals and the filter functions for that pulsar, as well as the inner product between the pairs of filter functions respectively, defined as
@@
where a filter function $\bm{\mathrm{S}}^{\mathrm{ki}}$ is a vector of size $\mathrm{N_{\rm TOA}}$ (which has the same structure as $\delta\bm{t}_{i}$).
@@
$\sigma_{i1}$ $=A\cos(\phi_{0}),$ (23)
$\sigma_{i2}$ $=-A\sin(\phi_{0}).$
@@
We can see from eq. (24) and eq. (25) that the number of shape and projection parameters in our model will go as $3\times N$ and $3+4\times N$ respectively for $N$ signal wavelets, since all wavelets share the same sky location and polarization angle. Notably, in previous work using the signal template defined by [23] for continuous waves, there are always the same number of shape and projection parameters. Our model has more projection than shape parameters, which will mean a larger percent of our model will be calculated with the more efficient methods in our code.
@@ II.3 Sampling
For Fisher matrix proposals, we can construct a matrix from the covariance between our model parameters. Then, the eigenvectors of the Fisher matrix can be used to change model parameters between MCMC steps based on the covariance between these model parameters. These Fisher matrices are constructed for various sets of model parameters separately from each other (i.e. for intrinsic pulsar noise, signal and noise transient wavelets, CURN, etc.), which allows us to make more informed MCMC jump proposals based on the type of step taken during sampling. Additionally, these fisher matrices can be recalculated during runtime based on the location of the MCMC sampler in parameter space to better inform our Fisher steps during sampling. For more details, see [29].
@@ II.4 Verification and Timing
Before analyzing efficiency improvements, we first present comparisons between the Bayesian likelihood implemented in QuickBurst and BayesHopperBurst. As BayesHopperBurst utilizes the software package enterprise [30] to evaluate the Bayesian likelihood, we use these values as a benchmark for our algorithm. The likelihood comparisons are evaluated using a simple data set consisting of two sine-Gaussian wavelets. Our comparisons show that the two signal models yield likelihoods that are equivalent up to $10^{-10}$ precision, as seen in Figure 1. Additionally, the signal reconstruction for a single pulsar from this run can be seen in Figure 2.
Figure 1: Likelihood differences between QuickBurst and enterprise corresponding to a simple dataset consisting of two sine-Gaussian signal wavelets. These do not exceed differences of $\sim 10^{-10}$ throughout the test. When including either intrinsic pulsar red noise or CURN, compounding numerical errors in likelihood evaluations increase differences to $\sim 10^{-5}$. However, the spread in likelihood differences is still centered around a difference of zero. The discrete steps in difference values are a result of numerical precision.
@@
With consistency between enterprise and QuickBurst established, we can now proceed to demonstrating the improved computational efficiency of QuickBurst. To do so, we simulate datasets as close to the 15 year NANOGrav dataset as possible (see §III for simulation details) and time our various types of steps. The timing for these steps correspond to runs on the SIM-MID dataset detailed in Table 2, which only vary our GW signal and noise transient parameters. However, in order to time noise jumps, we allow intrinsic pulsar RN to vary. It is worth noting that as the timing runs do not include CURN, which is handled in the ”Regular Fisher.” The average timing for these jumps is around 0.51 s if CURN is included. The results can be seen in Table 1.
@@
There are some notable takeaways from this timing data. First, we see QuickBurst yields a 1.5×2$\sim$ 1.5$\times$-2$\times$ 1.5 × - 2 × increase in time per step for jumps that vary shape parameters compared to enterprise. These include $\tau$-Scan, trans-dimensional (TD), and Regular Fisher jumps. Second, we see QuickBurst has an 8×8$\times$8 × increase in time per step compared to enterprise for Noise jumps, which vary intrinsic pulsar noise. This is primarily due to enterprise caching many of the inner products in calculating $C^{-1}$ in eq. (13), while QuickBurst re-constructs $C^{-1}$ during these steps, and then uses this for any future dot products.
@@ III.1 Datasets
for any signal h. For all three non-zero amplitude datasets, the signal SNR and luminosity distances can be found in Table 2. We calculate the SNR in two cases for each dataset: one case where we only include intrinsic pulsar WN in our covariance matrix, and another case where we include intrinsic pulsar RN and WN in our model. The difference between these SNR values indicates how much the detectability of the burst is effected by RN. The subsequent analysis performed on these four datasets in subsection III.2 includes both intrinsic pulsar RN and WN.
@@
Parameters $\mathrm{\tau(yrs)}$ $\mathrm{t_{0}(yrs)}$ $\mathrm{f_{0}(Hz)}$ $\mathrm{\theta_{\rm gw}}$ $\mathrm{\phi_{\rm gw}}$ $\mathrm{\psi_{\rm gw}}$ $\mathrm{A_{+}}$ $\mathrm{A_{\times}}$ $\mathrm{\phi_{0,+}}$ $\mathrm{\phi_{0,\times}}$
Range $0.2,5.0$ $0.0,T_{\rm max}$ $3.5e^{-9},1\times 10^{-7}$ $-\pi/2,\pi/2$ $0,2\pi$ $0,\pi$ $-10,-5$ $-10,-5$ $0,2\pi$ $0,2\pi$
@@ III.2 Analysis
Now we present the results of the analysis of datasets of varying signal strengths. We first analyze the no signal regime, and then analyze the datasets that have a signal present. To quantify our signal recovery, we can see how close the injected ($h$) and recovered ($h^{\prime}$) waveforms are to each other by computing a normalized noise weighted inner product between $h$ and $h^{\prime}$ over the entire PTA. This match $M$ is defined as
@@ III.2.2 Nonzero signal amplitude datasets
Next, for the recovered signal wavelets, it’s been seen in other generic GW burst algorithms that utilize a sine-Gaussian basis (see [36]), where the number of recovered wavelets scales with GW burst SNR, as the model can fit for more features in a higher SNR signal. This is consistent with the signal wavelet posteriors in Figure 4. Regarding the SIM-LOW dataset, we find that due to the significantly lower signal SNR, convergence took 3$\sim$ 3$\times$ 3 × longer than the SIM-MID and SIM-HIGH analyses. This is a result of marginal differences in the Bayesian likelihood when proposing GW signal wavelets at this SNR. Furthermore, we see there is a significant preference for no noise transient wavelets in all three cases, and that our reconstructions are always consistent with zero. Despite this, there is still a non-negligible percentage of samples at greater than zero noise transients. This effect is likely because the differences in the Bayesian likelihood when proposing adding a noise transient into our model being small. The model may occasionally try to add a noise transient, succeed, but proceed to remove it promptly in favor of excluding them in a later step.
@@
For the original SIM-MID search, the burst is only constrained to around a quadrant of the sky. We see, however, that including all NANOGrav 15 year pulsars (as shown in the middle figure) doesn’t provide additional constraints on our sky location recovery for this particular GW burst signal. The pulsars labeled by black dots are the additional pulsars we are adding, but there are no data for those pulsars around the burst epoch, despite lying in a more sensitive sky region.
To explore whether this is a feature of the SIM-MID dataset or our pipeline, we can simply change our sky location of the burst to have overlap with more of our longer timed pulsars. Moreover, we ensure that the total GW burst SNR is close to the signal SNR in the SIM-MID dataset. In doing so, we create a new dataset that is almost identical to the SIM-MID dataset, except the burst location is now at $\phi_{\mathrm{gw}}=5.0\ \mathrm{rad},\theta_{\mathrm{gw}}=\pi/2$, and $\psi_{\rm gw}=-0.3\ \mathrm{rad}$. Based on the bottom figure in Figure 5, it is clear that we very strongly recovery the injected burst location with a shifted burst location. In comparison with the sky location recovery efforts (as shown in the top and middle panels of Figure 5), we see that having better sky coverage allows us to better recover both $\theta_{\mathrm{gw}}$ and $\phi_{\mathrm{gw}}$, so these sky recovery differences are features of the datasets, and not a limitation of the sampler. We will discuss the issue of sky coverage and GW burst sensitivity more in Appendix Appendix A.
@@ III.2.3 Common red noise effects
The analysis we have discussed so far does not include a common uncorrelated red noise, or CURN. This shows the effectiveness of QuickBurst in a simpler context, but as we have seen from the NANOGrav 15 year GWB results [1], there is strong evidence for a correlated GWB. Therefore, we will show QuickBurst can disentangle a GW burst from a CURN process.
To isolate the effects a CURN process would have on our dataset, we test on a simple simulated data set of 20 pulsars with evenly sampled data that includes simulated pulsar WN, CURN, and a GW burst signal from a parabolic SMBH encounter at the same sky location described in the SIM-HIGH dataset (unlike the results in Figure 4, which used a more realistic dataset made to resemble the NANOGrav 15yr dataset). This dataset consists of a CURN injection with an amplitude of $A=6.0\times 10^{-15}$, a spectral index of $\gamma=13/3$, and $1\mu s$ of intrinsic pulsar WN (which sets the CURN process to be larger than the WN at the lowest frequencies).
@@
First, we perform a search with QuickBurst where we fix our CURN process to the injected CURN values. In doing so, we recover a significant portion of our injected signal, boasting a recovered SNR of $5.9$ (injected SNR of 6.8) and a match of $\mathrm{M=0.7}$. We then perform a run on the same dataset where we simultaneously vary both wavelet parameters and CURN spectral parameters, which is displayed in Figure 6 (a). Results indicate a match of $\mathrm{M=0.7}$ and a recovered SNR of 10.2 (injected SNR of 9.8). Furthermore, in comparison to histograms shown in Figure 4, the distribution of signal wavelets and noise transient wavelets are wider. This is likely due to covariance between the CURN process at overlapping frequencies, which can result in some of the.This suggests that even in strongly red noise dominated data, it is possible to disentangle the signal. There may be additional complications on real data in the future, where additional methods may help to better disentangle these two signal types, if a GW burst is present.

### v2

# ob A. Taylor0000-0001-9118-5589 d Burnette0009-0008-3649-0618 ce Bécsy0000-0003-0909-5563 l J. Cornish0000-0002-7435-0869

@@ II.1 Models
where $\bm{H}_{+}$ and $\bm{H}_{\times}$ are the plus and cross GW polarizations respectively, and have independent amplitudes $A_{+}$ and $A_{\times}$, with independent phases $\phi_{0,+}$ and $\phi_{0,\times}$. This leaves $t_{0},f_{0}$, and $\tau$ as common parameters between both polarizations. Furthermore, we model the phases independently, as not all GWs have the same 90 phase shift relationship between the plus and cross polarizations.
@@
where $F_{+}(\Omega_{k},\Omega_{\rm gw})$ and $F_{\times}(\Omega_{k},\Omega_{\rm gw})$ are the plus and cross polarization mode antenna responses, respectively. For further details on the geometric response, see [25]. Eq. (10) assumes that the only contribution to the signal is from the Earth term (ET) and not the Pulsar Term (PT). This is a result of our current observing baseline ($\sim 20$ years) being significantly shorter than the light travel time between the Earth and any pulsar in the PTA ($\sim 10^{3}$ years). Thus, any GW bursts which appear in only one pulsar will not affect the residuals in any other pulsar (unless two or more pulsars are within a few degrees of each other), nor will it reach the Earth within the time frame of the experiment. As such, this burst will appear as a single noise transient in that particular pulsar. For more details on the geometric time delay between the ET and the PT, see Section 2.1 in Ref. [7].
Then, for $N_{p}$ pulsars, the ET signal-to-noise ratio (SNR) is larger than the PT SNR by a factor of $\sqrt{N_{p}}$. The SNR scaling for the ET also becomes larger when considering that an ET search will allow us to see a common signal in all pulsars in our PTA, while the PT search will only show a signal in one pulsar and can be modeled as a noise transient. For additional details on these arguments, see [7, 26]. Moreover, it is expected that the number of events that we can possibly observe in the ET only search scales with the number of pulsars $\sim N_{p}^{3/2}$. However, the pulsar term (PT) only search scales as $\sim N_{p}$ due to the extended timing baseline when combining individual pulsar data. As a result, the number of possible observed events in the ET only search exceeds the number of possible observed events in the PT only search by $\sim N_{p}^{1/2}$. Considering current PTAs contain of order 100 pulsars, there’s potential to observe 10$\sim$ 10$\times$ 10 × more events in the ET only search.
@@ II.2 Likelihood
$\Sigma=B^{-1}+T^{T}N^{-1}T.\vspace{6pt}$ (14)
B is the prior matrix for the hyper-parameters, and T is the design matrix for the timing model, red noise (RN) and jitter (see [28]). The design matrix includes additional models for dispersion measure (DM), specifically modeling DM as an epoch-by-epoch DM offset. This model is refered to as DMX [28].
@@
$\log L=\sum_{i=1}^{N_{\mathrm{psr}}}\left[-\frac{1}{2}(\delta\bm{t}_{i}|\delta\bm{t}_{i})-\frac{1}{2}\log{\det(2\pi C_{i})}+\sum_{v=1}^{N_{F}}\big{[}\sum_{k=1,2}\sigma_{ki}N_{i}^{k}-\frac{1}{2}\sum_{k=1,2}\sum_{l=1,2}\sigma_{ki}\sigma_{li}M_{i}^{kl}\big{]}\right],$ (17)
where the last two terms are inner products between the $i^{th}$ pulsar residuals and the filter functions for that pulsar, as well as the inner product between the pairs of filter functions respectively. These inner products are defined as
@@
where a filter function $\bm{\mathrm{S}}^{\mathrm{ki}}$ is a vector of size $\mathrm{N_{\rm TOA}}$ (which has the same structure as $\delta\bm{t}_{i}$). Lastly, $N_{F}$ is the number of wavelets present in the signal model.
@@
$\sigma_{\textbf{1i}}$ $=A\cos(\phi_{0}),$ (23)
$\sigma_{\textbf{2i}}$ $=-A\sin(\phi_{0}).$
@@
We can see from eq. (24), the number of shape and projection parameters for noise transient wavelets in our model go as $3\times N_{n}$ and $2\times N_{n}$, respectively for $N_{n}$ noise transient wavelets. From eq. (25), the number of shape and projection parameters in our model will scale as $3\times N_{s}$ and $3+4\times N_{s}$ respectively for $N_{s}$ GW signal wavelets, since all wavelets share the same sky location and polarization angle. Notably, in previous work using the signal template defined by [23] for continuous waves, there are always the same number of shape and projection parameters. Our model has more projection than shape parameters, which will mean a larger percent of our model will be calculated with the more efficient methods in our algorithm.
@@ II.3 Sampling
For Fisher matrix proposals, we can numerically construct a matrix from the covariance between our model parameters by computing second derivatives of the likelihood via finite differences. Then, the eigenvectors of the Fisher matrix can be used to change model parameters between MCMC steps based on the covariance between these model parameters. These Fisher matrices are constructed for various sets of model parameters separately from each other (i.e. for intrinsic pulsar noise, signal and noise transient wavelets, CURN, etc.), which allows us to make more informed MCMC jump proposals based on the type of step taken during sampling. Additionally, these fisher matrices can be recalculated during runtime based on the location of the MCMC sampler in parameter space to better inform our Fisher steps during sampling. For more details on how these matrices are computed, see [29].
@@ II.4 Verification and Timing
Before analyzing efficiency improvements, we first present comparisons between the Bayesian likelihood implemented in QuickBurst and BayesHopperBurst. As BayesHopperBurst utilizes the software package enterprise [30] to evaluate the Bayesian likelihood, we use these values as a benchmark for our algorithm. The likelihood comparisons are evaluated using a simple data set consisting of two sine-Gaussian wavelets. Our comparisons show that the two signal models yield likelihoods that are equivalent up to $10^{-10}$ precision. When including either intrinsic pulsar red noise or CURN, compounding numerical errors in likelihood evaluations increase differences to $\sim 10^{-5}$. However, the spread in likelihood differences is still centered around a difference of zero. Additionally, the signal reconstruction for a single pulsar from this run can be seen in Figure 1.
@@
With consistency between enterprise and QuickBurst established, we can now proceed to demonstrating the improved computational efficiency of QuickBurst. To do so, we simulate datasets as close to the 15 year NANOGrav dataset as possible (see §III for simulation details) and time our various types of steps. The timing for these steps correspond to runs on the SIM-MID dataset detailed in Table 2, which only vary our GW signal and noise transient parameters. However, in order to time noise jumps, we allow intrinsic pulsar RN to vary. It is worth noting that the timing runs do not include CURN, which is handled in the ”Regular Fisher” jump proposals. The average timing for these jumps is around 0.51 s if CURN is included. The results can be seen in Table 1.
@@
There are some notable takeaways from this timing data. First, we see QuickBurst yields a 1.5×2$\sim$ 1.5$\times$-2$\times$ 1.5 × - 2 × increase in time per step for jumps that vary shape parameters compared to enterprise. These include $\tau$-Scan, trans-dimensional (TD), and Regular Fisher jumps. Second, we see QuickBurst has an 8×8$\times$8 × increase in time per step compared to enterprise for Noise jumps, which vary intrinsic pulsar noise. This is due to an increase in the number of inner products required for deterministic signal models. In expanding the likelihood to account for this deterministic model (as shown in eq. (17), we have a factor of $\frac{N_{F}^{2}+5N_{F}}{2}$ increase in the number of inner products we need to calculate during these Noise jumps in the M and N matrices seen in both eq. (18) and eq. (19), which accounts for the time increase shown.
@@ III.1 Datasets
for any signal h. For all three non-zero amplitude datasets, the signal SNR and luminosity distances can be found in Table 2. We calculate the SNR in two cases for each dataset: one case where we only include intrinsic pulsar WN in our covariance matrix, and another case where we include intrinsic pulsar RN and WN in our model. The difference between these SNR values indicates how much the detectability of the burst is affected by RN. The subsequent analysis performed on these four datasets in subsection III.2 includes both intrinsic pulsar RN and WN.
@@
Parameters $\mathrm{\tau(yrs)}$ $\mathrm{t_{0}(yrs)}$ $\mathrm{log_{10}(f_{0})}$ $\mathrm{\cos{\theta_{\rm gw}}}$ $\mathrm{\phi_{\rm gw}}$ $\mathrm{\psi_{\rm gw}}$ $\mathrm{A_{+}}$ $\mathrm{A_{\times}}$ $\mathrm{\phi_{0,+}}$ $\mathrm{\phi_{0,\times}}$
Range $0.2,5.0$ $0.0,T_{\rm max}$ $-8.46,-7$ $-1,1$ $0,2\pi$ $0,\pi$ $10^{-10},10^{-5}$ $10^{-10},10^{-5}$ $0,2\pi$ $0,2\pi$
@@ III.2 Analysis
Now we present the results of the analysis of datasets of varying signal strengths. We first analyze the no signal regime, and then analyze the datasets that have a signal present. Since the parameters of our generic model have no physical meaning, the pipeline’s performance is best characterized by how well the waveform of the GW signal is reconstructed. To quantify our signal recovery, we can see how close the injected ($h$) and recovered ($h^{\prime}$) waveforms are to each other by computing a normalized noise weighted inner product between $h$ and $h^{\prime}$ over the entire PTA. This match $M$ is defined as
@@ III.2.2 Nonzero signal amplitude datasets
Next, for the recovered signal wavelets, it’s been seen in other generic GW burst algorithms that utilize a sine-Gaussian basis (see [36]), where the number of recovered wavelets scales with GW burst SNR, as the model can fit for more features in a higher SNR signal. This is consistent with the signal wavelet posteriors in Figure 3. Regarding the SIM-LOW dataset, we find that due to the significantly lower signal SNR, convergence took 3$\sim$ 3$\times$ 3 × longer than the SIM-MID and SIM-HIGH analyses. This is a result of marginal differences in the Bayesian likelihood when proposing GW signal wavelets at this SNR. Furthermore, we see there is a significant preference for no noise transient wavelets in all three cases, and that our reconstructions are always consistent with zero. Despite this, there is still a non-negligible percentage of samples at greater than zero noise transients. Having no noise transients in our model is preferred, but having a few noise transients cannot be completely ruled out. This is expected, as white noise fluctuations can be fitted with a small amplitude wavelet.
@@
For the original SIM-MID search, the burst is only constrained to around a quadrant of the sky. We see, however, that including all NANOGrav 15 year pulsars (as shown in the middle figure) doesn’t provide additional constraints on our sky location recovery for this particular GW burst signal. The pulsars labeled by black dots are the additional pulsars we are adding, but there’s limited data for those pulsars around the burst epoch, despite lying in a more sensitive sky region. It’s also important to note that GW burst signals of this type have a finite width. While there are small changes in the posterior between the top two panels, the overall shape is the same between the 24 pulsar case and the 67 pulsar case. This is expected behavior given that these additional pulsars contribute little new GW burst information.
To explore whether this is a feature of the SIM-MID dataset or our pipeline, we can simply change our sky location of the burst to have overlap with more of our longer timed pulsars. Moreover, we ensure that the total GW burst SNR is close to the signal SNR in the SIM-MID dataset. In doing so, we create a new dataset that is almost identical to the SIM-MID dataset, except the burst location is now at $\phi_{\mathrm{gw}}=5.0\ \mathrm{rad},\theta_{\mathrm{gw}}=\pi/2$, and $\psi_{\rm gw}=-0.3\ \mathrm{rad}$. Based on the bottom figure in Figure 4, we clearly recover the injected burst location with a shifted burst location. In comparison with the sky location recovery efforts (as shown in the top and middle panels of Figure 4), we see that having better sky coverage allows us to better recover both $\theta_{\mathrm{gw}}$ and $\phi_{\mathrm{gw}}$, so these sky recovery differences are features of the datasets, and not a limitation of the sampler. We will discuss the issue of sky coverage and GW burst sensitivity more in Appendix Appendix A.
@@ III.2.3 Common red noise effects
The analysis we have discussed so far does not include a common uncorrelated red noise, or CURN. This shows the effectiveness of QuickBurst in a simpler context, but as we have seen from the NANOGrav 15 year GWB results [1], there is strong evidence for a correlated GWB. While it is beneficial to analyze the impacts of a correlated RN process on GW burst analysis, it is not expected that adding in Hellings-Downs (HD) correlations produces significant additional covariance with our wavelet model (see Appendix B for more details). Therefore, we will show QuickBurst can disentangle a GW burst from a CURN process.
To isolate the effects a CURN process would have on our dataset, we test on a simple simulated data set of 20 pulsars with evenly sampled data that includes simulated pulsar WN, CURN, and a GW burst signal from a parabolic SMBH encounter at a sky location of $\phi_{\mathrm{gw}}=3.7\ \mathrm{rad},\theta_{\mathrm{gw}}=\pi/2$, and $\psi_{\rm gw}=0.2\ \mathrm{rad}$ (unlike the results in Figure 3 which used a more realistic dataset made to resemble the NANOGrav 15yr dataset). This dataset consists of a CURN injection with an amplitude of $A=6.0\times 10^{-15}$, a spectral index of $\gamma=13/3$, and $1\mu s$ of intrinsic pulsar WN (which sets the CURN process to be larger than the WN at the lowest frequencies).
@@
First, we perform a search with QuickBurst where we fix our CURN process to the injected CURN values. In doing so, we recover a significant portion of our injected signal, boasting a recovered SNR of $6.4$ (injected SNR of 6.8) and a match of $\mathrm{M=0.82}$. We then perform a run on the same dataset where we simultaneously vary both wavelet parameters and CURN spectral parameters, which is displayed in Figure 5 (a). Results indicate a match of $\mathrm{M=0.79}$ and a recovered SNR of 10.2 (injected SNR of 9.0). Furthermore, in comparison to histograms shown in Figure 3, the distribution of signal wavelets and noise transient wavelets are wider. This is likely due to covariance between the CURN process at overlapping frequencies, which can result in some of the GW burst power being absorbed into the CURN model. This suggests that even in strongly red noise dominated data, it is possible to disentangle the signal. There may be additional complications on real data in the future, where additional methods may help to better disentangle these two signal types, if a GW burst is present.